Sarah is 5 times older than her grandaughter. If the difference between their ages is 40, how old is her granddaughter?
Answer by dabanfield(803)

Sarah is 5 times older than her grandaughter. If the difference between their ages is 40, how old is her granddaughter?

Let x be the grandaughter's age and y be Sarah's age. Then from the two statements we know that:
1.) y = 5*x
2.) y - x = 40
Substituting 5x for y in equation 2.):
5x - x = 40
4x = 40
x = 10 (the grandaughter's age)
Calculate Sarah's age from 1.) as 5*x = 5*10 = 50